adult tickets for a play cost $5 and children tickets cost $2 for 1 person. 460 tickets were sold. Receipts were $1880. Find # of adult tickets sold. Please help me with this problem!
Answer by tutorcecilia(2144)

(Rate of A)(Amount of A) + (Rate of B)(Amount of B) = Amount of C
$5.00(x)+$2.00(460-x)=$1880 [x+y=460, so 460-x=y)]
5x+2(460-x)=1880 [Simplilfy]
5x+920-2x=1880 [Solve for x]
5x-2x=-920+1880
3x=960
x=320 adult tickets sold
.
Since x+y=460 and 460-x=y, than 460-320=140 children tickets sold
Checking: 320+140=460
.
Checking:
5x+2(460-x)=1880 [Plug-in the values for (x=320) and (y=140)]
5(320)+2(460-(320))=1880
1880=1880 [Checks out]